OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (9,3); slope:4/9 y – 9 = (x + 3) y – 9 = (x – 3) y + 3 = (x – 9) y – 3 = (x – 9)

OpenStudy (maeganr):

y-y1=m(x-x1) is the formula for point-slope form.

OpenStudy (maeganr):

So, your points are (9,3) and your slope is 4/9.

OpenStudy (maeganr):

9:x 3:y y-3=4/9(x-9) Is the equation. Now do you have to solve it also?

OpenStudy (maeganr):

Hello? Are you there? O.o

OpenStudy (anonymous):

so then it would be D?

OpenStudy (maeganr):

Yes. but it doesn't have the slope=m in it. Make sure the answer is basically the same exact thing as y-3=4/9(x-9) .

OpenStudy (maeganr):

does*